OpenStudy (anonymous):

Which of the following equations describes the line shown below? Check all that apply. A. y - 4 = -2(x + 4) B. y - 1 = -(x - 2) C. y - 2 = -(x - 1) D. y - 4 = -(x - 4) E. y - 1 = -2(x - 2) F. y - 4 = -(x + 4)

OpenStudy (precal):

Were you given a graph?

OpenStudy (anonymous):

no

OpenStudy (anonymous):

http://media.apexlearning.com/Images/200706/18/934e2b10-a637-4380-abd7-96b23a05b97b.gif A. y - 6 = (x - 12) B. y - 6 = 4(x - 12) C. y - 12 = (x - 6) D. y - 4 = (x - 4) E. y - 12= 4(x - 6) F. y - 4 = 4(x - 4) help pleas!

OpenStudy (precal):

sorry not able to view link at all.